RodeWolf:

It is 76º F at the 6000-foot level of a mountain, and 49º F at the 12,000-foot level of the mountain. Write a linear equation, in slope-intercept form, to find the temperature T at an elevation e on the mountain, where e is in thousands of feet.

jhonyy9:

do you know what is the slope intercept form ?

RodeWolf:

No

swinn:

Find the slope using the slope formula: m = %28y2-y1%29%2F%28x2-x1%29 Assign the given values as follows x1 = 6 (for 6000 ft); y1 = 76 x2 = 12(for 12000ft); y2 = 49 m = 49-7/12-6 = -27/6 = -9/2 Use the point slope equation y - y1 = m(x - x1) so create the equation y - 76 = -9/2(x - 6) y - 76 = -9/2x - (-27) y - 76 = -9/2x + 27 + 76 y = -9/2x + 103; is the equation Using e and T T = -9/2e + 103 https://www.algebra.com/algebra/homework/coordinate/word/Linear_Equations_And_Systems_Word_Problems.faq.question.185353.html